A toy rocket is fired at v0 = 40.7 m/s at an angle of ? = 72.3o with respect to the horizontal on flat level ground. Assuming that air resistance is negligible, what is its maximum height (H) and how far (R) will the rocket have traveled in the horizontal direction by the time it lands?
Answer #1

0.5*M*Vy^2 = M*g*H

H = Vy^2/2g = (V*sin72.3)^2/2g = (40.7*sin72.3)^2/(2*9.81) =76.62m

R = V^2*sin(2*72.3)/g = (40.7^2*sin(72.3*2))/9.8 = 97.92 m
Answer #2

v=40.7

H=(v*sintheta)^2/(2*g)

theta=72.3

g=9.8m/s^2

H=76.7m

R=v^2*sin(2*theta)/(2*g)

=48.95m
Answer #3

given: U=40.7m/s

?=72.30